An antelopes moving with constant acceleration covers the distance between two points 70.0m apart in 7.0s. It’s speed aa it passes the second point is 15.0m/s


A.) What is it’s speed at the first point? b) What is the acceleration?



1
Expert's answer
2021-03-29T09:03:59-0400

(a)

"d=\dfrac{1}{2}(v_i+v_f)t,""v_i=\dfrac{2d}{t}-v_f,""v_i=\dfrac{2\cdot70\ m}{7\ s}-15\ \dfrac{m}{s}=5\ \dfrac{m}{s}."

(b)

"v_f=v_i+at,""a=\dfrac{v_f-v_i}{t}=\dfrac{15\ \dfrac{m}{s}-5\ \dfrac{m}{s}}{7\ s}=1.43\ \dfrac{m}{s^2}."